OpenStudy (anonymous):

Write the polynomial of the smallest degree with roots i, -i, 4i, and -4i.

jimthompson5910 (jim_thompson5910):

x = i, x = -i, x = 4i, or x = -4i x = +-i or x = +-4i x^2 = -1 or x^2 = -16 x^2 +1 = 0 or x^2+16 = 0 (x^2+1)(x^2+16)=0 I'll let you finish it up